OpenStudy (anonymous):

A 520 m wire was cut into three pieces. The second piece is 4 times as long as the first. The third piece is 2 times as long as the second. How long is each piece?

OpenStudy (anonymous):

x +4x+2x=520. That's what I get and then I get stuck.

OpenStudy (anonymous):

if the first piece has length x, then the second piece has length 4x, and the third piece has length 8x (it is twice as long as the SECOND piece). So x + 4x + 8x = 520. Add up the x's and you get 13x = 520. Divide both sides by 13 and you get your answer.

OpenStudy (anonymous):

I get x + 4x + 8x = 520. Combine like terms to get 13x = 520. Divide both sides to get x = 40.